The magnetic field in a region is given by B = B0( 1 + x/a)k. A square loop of edge length d is placed with its edge along x and y axes. The loop is now moved with constant velocity V = V0i.The emf induced in the loop at that instant is

Grade:12

2 Answers

parthasarathy
9 years ago
Emf=B*l*v =Bo(1+x/a)*d*Vo
Himaja
9 years ago
emf = d(flux)/dt
= d(BA)/dt
= Ad(1+x/a)B0/dt
=d2B0V0/a